### v3

@@ III Simulation Results and Analysis
Figure 3: Mean return across clean and visibility-noise environments. PPO+LLM performs comparably in clean settings and maintains advantage under perceptual uncertainty.
Figure 4: Mean return across noisy environments (Order Timing Noise and Combined Noise). PPO+LLM demonstrates improved robustness over PPO under multi-agent coordination disruptions. Results averaged over n=200 evaluation episodes.
We conduct a systematic evaluation of our PPO+LLM framework using the Overcooked-AI environment to analyze how LLM-guided reward shaping influences multi-agent collaboration, task efficiency, and robustness to uncertainty. Each model is tested over 200 independent episodes per environment, with controlled random seeds and standardized starting conditions to support consistent and fair evaluation.
We evaluate both standard PPO and PPO+LLM agents across four environments of increasing complexity: Clean (no noise), Visibility Noise, Order Timing Noise, and Combined Noise. The metrics include returns per episode, standard deviation, and average per-step latency. We report and analyze results across three core dimensions: performance, robustness, and inference efficiency.
Figure 5: Qualitative comparison in Overcooked-AI. Left: PPO+LLM agents coordinate by specializing in distinct roles and maintaining clear handoffs. Right: PPO-only agents often block each other, duplicate roles, or stall at delivery. The frozen LLM encourages efficient collaboration by rewarding high-level intent and coordination. Results averaged over n=200 evaluation episodes.
III-A Returns per Episode
Figure 3 summarizes mean rewards per episode for PPO and PPO+LLM. In the Clean setting, where agents observe the environment without any disturbance, PPO fails to achieve any reward. This indicates that the agent is unable to develop cooperative strategies without step-by-step feedback guiding its actions during the episode. In contrast, PPO+LLM consistently achieves a reward of 9.8 with zero variance, demonstrating that binary LLM evaluations are sufficient to support reliable task completion. Under the Visibility Noise condition, where the agent’s input state is interrupted with a random amount of perceptual noise, PPO again collapses to a return of 0. This outcome shows PPO’s sensitivity to low-level visual disturbances. In contrast, PPO+LLM stays strong and achieves a perfect return of 9.8 again. Since it learns from prompt-driven evaluations that do not rely on precise raw data, PPO+LLM remains stable even when perception is inconsistent.
Figure 4 shows the average return per episode under structured disturbances, meaning controlled and repeatable disruptions to either perception (observation noise) or task flow (order timing noise). In the Order Timing Noise condition, where task arrivals are randomly delayed or reordered, PPO performs poorly and has an average return of 40.4-40.4 (standard deviation 3.7). Negative returns occur because agents accumulate time-step penalties and failed delivery costs without successfully completing enough tasks to balance these losses, meaning policies waste time or act inefficiently such that penalties outweigh rewards. PPO+LLM significantly mitigates this degradation, recovering over 10 points (mean: 30.3-30.3, std: 4.4) by consistently evaluating overall task progression (e.g., timestep and pending orders), which reduces instability caused by noisy timing dynamics. Under Visibility Noise, where the grid-level tensor is partially corrupted before reaching the policy network, PPO again collapses due to its reliance on unaltered input encodings for fine-grained value learning, resulting in no return accumulation. In contrast, PPO+LLM maintains positive returns by relying on symbolic evaluations that remain stable even when perception is degraded. Together, these results demonstrate that PPO+LLM is substantially more resilient than PPO under both timing and observation disturbances, highlighting the importance of symbolic feedback for robust coordination in noisy multi-agent environments.
This pattern stays the same in the Combined Noise condition, where both visibility and timing are altered. PPO deteriorates to a mean of 39.9-39.9 (std: 3.6), while PPO+LLM recovers nearly 25% of that loss with a mean return of 29.6-29.6 (std: 4.3). This shows how symbolic feedback helps learning when rewards are weak or inconsistent, by giving the agent a clearer understanding of progress. To illustrate how the system works in practice, we present a qualitative example from the Overcooked-AI environment (Figure 5). In this episode, one agent retrieves ingredients while the other prepares plates and manages deliveries. The frozen LLM rewards states where agents remain active, coordinate handoffs efficiently, and avoid interfering with one another. This leads to smooth, high-throughput cooperation. In contrast, the PPO-only baseline often results in role duplication, hesitation, or spatial conflicts between agents. This example shows that symbolic feedback improves not just numerical performance but also the structure and clarity of collaborative behavior.
Figure 7: Burger Kitchen: Coordination Under Observation Noise. Left: PPO+LLM agents maintain effective collaboration despite partial visual occlusion. Even when perceptual details are masked or distorted, symbolic feedback helps agents preserve task flow and avoid conflicts. Right: PPO baseline agents fail to recognize delivery opportunities due to limited visibility, leading to idle burgers and missed outputs. This figure demonstrates how language-based shaping enables perceptual robustness and consistent goal pursuit under noisy observations.
Figure 8: Burger Kitchen: Coordination Under Combined Noise. Left: PPO+LLM agents handle both mis-timed ingredient requests (clock icon) and sensor corruption (pixel-noise tile), reassigning roles on the fly and completing delivery. Right: PPO baseline agents break the task pipeline; perception noise triggers hesitation, timing jitter causes the burger to arrive too late, and the meal sits idle ($\times$) at the window. This comparison shows how prompt-based symbolic feedback enables recovery from multiple simultaneous disturbances.
To evaluate coordination performance under temporal uncertainty, we test both PPO and PPO+LLM agents in the Order Timing Noise environment, where requested ingredients arrive at irregular intervals. This setting introduces misaligned task dependencies, increasing the risk of agents acting prematurely or stalling due to incomplete orders. As shown in Figure 6, PPO+LLM agents recover from off-cycle prompts (e.g., a tomato requested ahead of time) by dynamically reassigning roles and maintaining structured task flow. In contrast, PPO baseline agents lack a mechanism to reconcile timing inconsistencies, leading to delivery failures and idle outputs. This scenario highlights the effectiveness of prompt-based symbolic guidance in helping agents reason through dynamic task adjustments, enabling more stable coordination even when order streams are perturbed.
To evaluate robustness under perceptual uncertainty, we test agents in the Observation Noise environment, where input states are partially masked or corrupted prior to reaching the policy network. This simulates scenarios in which agents must act with incomplete or degraded visibility. As shown in Figure 7, PPO+LLM agents successfully maintain cooperative behavior, relying on symbolic evaluations that remain stable across noisy visual inputs. The frozen LLM reinforces high-level intent rather than low-level pixel accuracy, enabling agents to avoid collisions, infer teammate roles, and complete deliveries even when specific object placements are obscured. In contrast, PPO baseline agents often misinterpret the scene, leading to duplicated efforts or missed handoffs. This setting highlights the value of prompt-based feedback in providing consistent behavioral signals, especially when raw observations are unreliable.
To evaluate agent performance under simultaneous disturbances, we consider the Combined Noise environment, which introduces both observation corruption and task timing irregularities. This dual-source disruption challenges agents to interpret partially noisy visual inputs while also adjusting to asynchronous order arrivals. As shown in Figure 8, PPO baseline agents frequently fail in this setting. Visual distortion causes hesitation or miscoordination, and mistimed ingredient requests result in broken task flow, leading to idle or undelivered meals. In contrast, PPO+LLM agents use symbolic feedback to remain adaptive. The LLM reinforces cooperative behavior even under perceptual ambiguity, enabling agents to reassess roles and recover from misalignments. This robustness allows PPO+LLM to maintain productive collaboration despite multiple overlapping sources of uncertainty, demonstrating the effectiveness of language-based shaping in complex, real-time environments.
III-B Robustness to Noise
Since PPO+LLM’s prompt-based evaluation relies on higher-level elements like order count and timestep, which are not directly harmed by visual corruption, it remains unaffected and keeps an average return of 9.8.
Order Timing Noise introduces uncertainty into the environment’s causal structure by changing when new objectives are presented. This disrupts the routine coordination patterns PPO often overfits to. Without the ability to anticipate task sequences, PPO’s policy collapses. PPO+LLM, however, keeps its strategic framework intact by relying on ongoing LLM-guided assessments to judge whether progress is being made. Its behavior remains adaptive, shaped by a reward signal that is context-aware and flexible.
Combined Noise, which is arguably the most difficult condition, introduces both perceptual distortion and goal-timing irregularity, disrupting low-level coordination patterns. PPO’s performance deteriorates substantially in this setting. In contrast, PPO+LLM continues to produce reliably higher returns by leveraging symbolic feedback from LLM prompts. As shown in Figure 4, this feedback serves as an external corrective mechanism, helping agents focus on high-level cooperation rather than brittle stimulus-response loops. This enables more resilient coordination even when the environment becomes unpredictable or misaligned with learned routines.
Figure 9: Latency per decision step across environments. PPO+LLM adds modest computational cost but stays within real-time limits across all conditions. Results averaged over n=200 evaluation episodes.
@@ III-C Latency Tradeoff
Figure 9 summarizes the latency per decision step across four environments. While PPO+LLM delivers improvements in robustness and strategic coordination, it introduces a lightweight runtime cost due to real-time prompt generation and LLM-guided evaluation. In the Clean environment, PPO’s per-step latency averages 0.77 milliseconds, whereas PPO+LLM increases this to 0.99 milliseconds, introducing an approximately 28% overhead caused by token inference and string formatting. Similar latency increases are observed across all settings, but PPO+LLM consistently remains below 1.05 milliseconds per step, ensuring that performance remains within real-time constraints.
Importantly, this overhead is both small and stable. Despite task randomness and structured disturbances, the shaping module’s runtime does not fluctuate due to its reliance on a frozen LLM and a deterministic prompt structure. This predictability is critical for downstream integration in safety-sensitive or latency-bound applications. In addition, because the LLM is queried with compact, template-based prompts, the shaping mechanism remains computationally efficient and scales well across multi-agent deployments. These properties make PPO+LLM viable not just for simulated benchmarks, but also for interactive robotics, collaborative agents, and fast-paced games where decision-time guarantees are essential.
Figure 10: Task-Completion Rate ($\geq 9.4$) across environments. PPO+LLM reliably completes full tasks in clean and perceptually noisy settings, while PPO consistently fails to reach the success threshold. Results averaged over n=200 evaluation episodes.
III-D Task Completion Rate
To evaluate whether agents successfully completed the full collaborative task, we introduced a completion-rate metric that counts the percentage of episodes where the total return exceeded a threshold of 9.4. As shown in Figure 10, this cutoff was chosen based on empirical observations, where PPO+LLM agents consistently reached scores near 9.8 in successful episodes. Setting the threshold slightly below 9.8 allows for small variations while still representing full task completion.
As shown in Figure 10, PPO+LLM agents achieved a 100% completion rate in both the clean and observation noise settings, while PPO agents consistently failed to meet the threshold in all environments. In more challenging conditions, such as order timing noise and combined noise, neither policy consistently completed the task, resulting in 0% success. However, PPO+LLM still demonstrated substantially higher average returns in these noisy environments, indicating partial progress toward task completion despite increased difficulty.
These results suggest that LLM-guided agents are more capable of executing full task cycles in favorable or moderately noisy environments and maintain more robust performance under degraded conditions, even when full completion is not achieved. To qualitatively demonstrate this coordination, next we visualize the behavioral progression toward equilibrium in a concrete kitchen setting.
Figure 11: Convergence to Nash-like equilibrium in a 2-agent Burger Kitchen task. PPO-LLM agents transition from initial miscoordination (Step 1) to stable cooperative behavior (Step 4), resulting in a joint policy resembling equilibrium stability, where neither agent benefits from deviating unilaterally.

### v4

@@ III Simulation Results and Analysis
We now present a quantitative evaluation of the PPO+LLM framework in the two–agent Overcooked Markov game described in Section II. The goal of this section is to connect the implementation described in the training and analysis scripts with measurable performance, robustness, equilibrium behavior, and runtime characteristics.
Unless otherwise noted, all experiments use the asymmetric_advantages layout with horizon $H=400$, the four environment regimes
$\mathcal{E}=\{\text{No Noise},\;\text{Noise},\;\text{Delay},\;\text{Combo}\},$
and the six baselines
$\mathcal{B}=\{$ $\text{Baseline},\;\text{PPO+LLM},\;\text{CC\_PPO},$
$\text{SP\_PPO},\;\text{HARL},\;\text{PBT\_PPO}\}.$
For each $(m,e)\in\mathcal{B}\times\mathcal{E}$ we train five random seeds
$\mathcal{S}_{\text{train}}=\{1001,\;2002,\;3003,\;4004,\;5005\}$
using the evaluation procedure described in Section II. Baseline, CC_PPO, SP_PPO, HARL, and PBT_PPO are trained for $10^{6}$ environment steps, while PPO+LLM is trained for $6\times 10^{5}$ steps. After training, each triple $(m,e,r)$ is evaluated for $N_{\text{eval}}=10$ episodes, producing a mean episodic return and empirical standard deviation. All aggregate statistics reported below are averages over these five seeds unless otherwise stated, and are computed directly from the outputs of the evaluation pipeline.
Whenever possible we report both (i) seed averaged summary statistics in tables and (ii) box–and–swarm plots that visualize all $N=5$ seeds. These distributions make it clear that most differences between baselines lie in a narrow performance band and that apparent gaps are rarely statistically significant.
III-A Episodic Return Across Baselines
For each model we denote by $\hat{V}^{m,e,r}$ the empirical mean episodic return in environment $e$ for baseline $m$ and seed $r$, computed over $N_{\text{eval}}$ evaluation episodes as
$\hat{V}^{m,e,r}=\frac{1}{N_{\text{eval}}}\sum_{i=1}^{N_{\text{eval}}}G^{(i)},$ (18)
where $G^{(i)}$ is the return of evaluation episode $i$ as in (32). We summarize performance across seeds by the seed averaged value
$\bar{V}^{m,e}=\frac{1}{|\mathcal{S}_{\text{train}}|}\sum_{r\in\mathcal{S}_{\text{train}}}\hat{V}^{m,e,r},$ (19)
and use the empirical standard deviation over seeds to indicate sensitivity to random initialization.
Across all baselines, the No Noise and Noise regimes are dominated by sparse rewards. Empirically every baseline satisfies
[…]
with zero empirical standard deviation across seeds in both regimes. Under the current horizon and reward configuration, policies almost never obtain positive task level reward in these regimes and instead hover exactly at the zero baseline. As a result, No Noise and Noise act primarily as sanity checks on the training script and architecture rather than as discriminative stress tests.
The Delay and Combo regimes are more informative. Recall from Section II that the Delay wrapper subtracts a penalty of $0.5$ with probability $0.2$ at every timestep. Over a horizon of $H=400$ steps the expected cumulative penalty is
$\mathbb{E}\Bigg[\sum_{t=0}^{H-1}r_{t}^{\text{delay}}\Bigg]\approx-0.1\times 400=-40,$
even if the agents never complete a single order. The empirical means confirm that all methods operate very close to this theoretical penalty floor. Aggregating over seeds, Table I summarizes $\bar{V}^{m,e}$ and its seed standard deviation for the penalty regimes.
TABLE I: Seed averaged episodic return $\bar{V}^{m,e}$ in the penalty regimes. Values are mean $\pm$ standard deviation across $5$ seeds.
Baseline Delay Combo
Baseline $38.74\pm 0.96$ $40.25\pm 0.79$
PPO+LLM $39.51\pm 1.94$ $39.19\pm 0.88$
CC_PPO $38.74\pm 0.96$ $40.25\pm 0.79$
SP_PPO $40.13\pm 1.55$ $39.45\pm 1.80$
HARL $40.18\pm 1.38$ $40.22\pm 0.81$
PBT_PPO $38.79\pm 1.00$ $39.17\pm 0.51$
Figure 7: Episodic return distributions across clean and perturbed regimes (N = 5 seeds per configuration).
For the Delay regime we obtain
TABLE II: Seed-averaged episodic return in the Delay regime. Values are mean $\pm$ standard deviation across $5$ seeds.
Baseline Delay
Baseline $38.74\pm 0.96$
CC_PPO $38.74\pm 0.96$
PBT_PPO $38.79\pm 1.00$
PPO+LLM $39.51\pm 1.94$
SP_PPO $40.13\pm 1.55$
HARL $40.18\pm 1.38$
All values lie within roughly $1.3$ reward units of the theoretical floor at 40-40. Baseline and CC_PPO achieve the highest (least negative) mean return at 38.74-38.74, PBT_PPO is essentially tied at 38.79-38.79, while PPO+LLM, SP_PPO, and HARL sit slightly below, between 39.5-39.5 and 40.2-40.2. The seed standard deviations in Delay range from about $0.96$ to $1.94$, indicating moderate but not extreme sensitivity to random initialization.
For the Combo regime, which combines the same delay penalty with additive Gaussian observation noise, we obtain
TABLE III: Seed-averaged episodic return in the Combo regime. Values are mean $\pm$ standard deviation across $5$ seeds.
Baseline Combo
PBT_PPO $39.17\pm 0.51$
PPO+LLM $39.19\pm 0.88$
SP_PPO $39.45\pm 1.80$
HARL $40.22\pm 0.81$
Baseline $40.25\pm 0.79$
CC_PPO $40.25\pm 0.79$
Here the average return across all baselines is approximately 39.76-39.76, again within a single reward unit of the 40-40 penalty floor. PBT_PPO and PPO+LLM separate slightly from the remaining baselines, recovering about one reward unit relative to Baseline and CC_PPO. The gap between PPO+LLM and the raw PPO Baseline in Combo is
$\bar{V}^{\text{PPO+LLM},\text{Combo}}-\bar{V}^{\text{Baseline},\text{Combo}}\approx 1.06,$
which corresponds to about $2.6\%$ of the magnitude of the penalty floor. Because each unit of return accumulates over many timesteps, this seemingly small difference reflects a nontrivial reduction in unnecessary penalties and/or an increase in successful partial task completions.
A one–way ANOVA over the $N=5$ seeds for each penalty regime does not detect statistically significant differences between baselines. In Delay we obtain $F(5,24)\approx 1.31$ and $p\approx 0.29$, and in Combo $F(5,24)\approx 1.42$ and $p\approx 0.25$. The heavy overlap of boxes and points in Fig. 7(b) makes this lack of clear separation visually apparent. Overall, PPO+LLM sits in the middle of this family. In Delay it is slightly worse than the raw PPO Baseline but still close to the penalty floor; in Combo it is essentially tied with PBT_PPO for the best performance and improves notably over the Baseline, CC_PPO, and HARL.
III-B Robustness to Noise and Structured Penalties
To study robustness, we examine how returns shift as we activate different environment wrappers. For each baseline $m$ we define the raw performance drop relative to the clean regime,
$\begin{split}\Delta_{\text{noise}}^{m}&=\bar{V}^{m,\text{Noise}}-\bar{V}^{m,\text{No Noise}},\\ \Delta_{\text{delay}}^{m}&=\bar{V}^{m,\text{Delay}}-\bar{V}^{m,\text{No Noise}},\\ \Delta_{\text{combo}}^{m}&=\bar{V}^{m,\text{Combo}}-\bar{V}^{m,\text{No Noise}}.\end{split}$ (20)
Because all methods satisfy $\bar{V}^{m,\text{No Noise}}=0$, these quantities coincide with the raw means in the perturbed regimes. The robustness script also records the absolute magnitude of these drops as
$D^{m}_{\text{delay}}$ $=-\Delta_{\text{delay}}^{m},$
$D^{m}_{\text{combo}}$ $=-\Delta_{\text{combo}}^{m},$
so that a larger $D$ corresponds to a larger loss relative to the clean regime.
Figure 8: Distribution of robustness drops $D^{m}_{\text{delay}}$ and $D^{m}_{\text{combo}}$ relative to the No Noise regime (N = 5 seeds). Lower values indicate smaller losses. All baselines cluster tightly around the theoretical $40$ point penalty.
Across baselines we observe that
$\Delta_{\text{noise}}^{m}=0,\qquad D^{m}_{\text{noise}}=0\quad\forall\,m\in\mathcal{B}.$
Pure observation noise, implemented as additive Gaussian perturbations to the feature vector, has almost no measurable effect in the current training budget. Policies have not learned fine grained strategies that depend sensitively on the featurization, so perturbing observations without changing the reward structure does not change the final returns.
In contrast, the Delay and Combo regimes induce large and highly structured drops in value. The robustness aggregation script reports, for every $m\in\mathcal{B}$,
$D^{m}_{\text{delay}}\in[38.74,40.18],\qquad D^{m}_{\text{combo}}\in[39.17,40.25],$
with seed standard deviations on the order of one reward unit. These values are in tight agreement with the theoretical $40$ point penalty derived above, and show that nearly all of the return in the perturbed regimes is accounted for by the known delay penalty rather than by frequent catastrophic failures. Figure 8 makes this visually explicit: each baseline occupies a narrow vertical band just below forty.
Within this frame, PPO+LLM exhibits two meaningful robustness properties.
  1. 1.
    In the Delay regime, PPO+LLM does not collapse catastrophically: its drop $D^{\text{PPO+LLM}}_{\text{delay}}\approx 39.51$ is comparable to the drops of Baseline and CC_PPO ($38.74$) and lies comfortably inside the range observed for all methods. The symbolic shaping signal neither amplifies the effect of penalties nor destabilizes training in this setting, which is an important safety property when rewards are heavily perturbed.
  2. 2.
    In the Combo regime, PPO+LLM slightly reduces the impact of combined observation noise and delay relative to the raw Baseline. Baseline and CC_PPO realize drops of $D^{\text{Baseline}}_{\text{combo}}=D^{\text{CC\_PPO}}_{\text{combo}}=40.25$, while PPO+LLM reduces this to $D^{\text{PPO+LLM}}_{\text{combo}}\approx 39.19$, a recovery of about $1.06$ reward units. When we compare models seed by seed, PPO+LLM achieves a higher mean return than the corresponding Baseline model in $4$ out of $5$ seeds in this regime. This suggests that the shaping bonus helps the policy maintain a consistent sense of cooperative progress when both perception and timing are perturbed.
A one–way ANOVA on $D^{m}_{\text{delay}}$ and $D^{m}_{\text{combo}}$ across baselines again finds no statistically significant differences (Delay: $F(5,24)\approx 1.31$, $p\approx 0.29$; Combo: $F(5,24)\approx 1.42$, $p\approx 0.25$), which is consistent with the highly overlapping distributions in Figure 8. The remaining baselines further contextualize these trends. PBT_PPO and SP_PPO show modest gains over the Baseline in Delay, indicating that exploration schemes and self play can help agents adapt their timing to structured penalties. HARL, which uses a simple hand designed shaping term based on remaining orders, demonstrates the risk of naive shaping: when the shaping signal is not carefully calibrated, it can amplify the effect of penalties rather than compensate for them. In contrast, PPO+LLM uses a bounded shaping coefficient and a frozen LLM to provide stable, context dependent feedback that respects the original reward scale.
To give qualitative intuition, we also visualize coordination in the original dense reward Burger Kitchen layout described in Section II. Figure 2 shows the clean task layout, and Figures 36 then show PPO+LLM and PPO baselines in the presence of order timing noise, observation noise, and combined disturbances. Across all three settings, PPO+LLM maintains coherent task flow while PPO frequently stalls or miscoordinates, mirroring the robustness trends observed in the asymmetric_advantages layout.
@@ III-C Latency Tradeoff
Beyond reward, practical deployment requires that shaping respect real time constraints. To quantify runtime cost, we instrument the environment wrappers to record wall clock time per decision step. For each trained policy we run a separate evaluation phase without gradient updates, execute $5$ full episodes in each regime, and measure the elapsed time required to generate actions and apply any shaping logic. We then compute a per step latency in milliseconds and aggregate across seeds in the same way as for return. Table IV reports the seed averaged per step latency for Baseline PPO and PPO+LLM in each regime.
Figure 9: Inference latency distribution (N = 5 seeds). Each box aggregates all regimes for a given baseline. All methods lie in a narrow band around $1.6$ ms per decision step.
For the raw PPO Baseline, the mean per-step inference latency averaged over all four environment regimes and all seeds is approximately $1.60\,\mathrm{ms}$. Table IV reports the regime-specific latency statistics. Across regimes, inference latency varies by less than $0.03\,\mathrm{ms}$, indicating that neither observation noise nor structured delay penalties introduce meaningful runtime overhead.
PPO+LLM remains in the same latency range. Despite the additional symbolic evaluation step, its per-step latency is statistically indistinguishable from the Baseline across all regimes, with an overall average of approximately $1.60\,\mathrm{ms}$.
TABLE IV: Mean per-step inference latency (ms) across environment regimes. Values are mean $\pm$ standard deviation across $5$ seeds.
Method No Noise Noise Delay Combo
Baseline $1.59\pm 0.01$ $1.62\pm 0.02$ $1.59\pm 0.00$ $1.61\pm 0.01$
PPO+LLM $1.58\pm 0.00$ $1.60\pm 0.01$ $1.60\pm 0.01$ $1.63\pm 0.02$
The worst case overhead relative to the Baseline occurs in the Combo regime, where PPO+LLM is slower by about $1.36\%$:
$\frac{1.634-1.612}{1.612}\approx 1.4\%.$
In the other regimes PPO+LLM is within $1\%$ of the Baseline, and in No Noise the measured mean is slightly lower due to random variation. A global one–way ANOVA on latency across all baselines yields $F(5,174)\approx 0.43$ and $p\approx 0.83$, confirming that latency differences between methods are not statistically significant. Figure 9 shows that all baselines occupy essentially the same 1.581.581.641.64 ms range.
Training time tells a complementary story. Using the same Colab GPU backend, the Baseline requires approximately $520$ minutes to complete a full $10^{6}$ step training run
$T_{\text{Baseline}}\approx 519.7\pm 0.3\,\text{minutes},$
while PPO+LLM, trained for $6\times 10^{5}$ steps, completes in
$T_{\text{PPO+LLM}}\approx 295.3\,\text{minutes (median)},$
with a mean of $285.3$ minutes across seeds and some early terminated runs that finish even faster. Relative to the Baseline, PPO+LLM reduces the median wall clock training time by about $43\%$:
$1-\frac{295.3}{519.7}\approx 0.432.$
Taken together, these results show that PPO+LLM trades a small, roughly one percent per step inference overhead for a significantly shorter training schedule and modest robustness gains in the most difficult regimes, while remaining well within typical real time control budgets.
III-D Task Completion Under Sparse Rewards
In earlier Burger Kitchen experiments with dense task rewards it was natural to measure task completion by counting episodes whose return exceeded a threshold close to the optimal score (for example, a threshold of $9.4$ when successful episodes consistently achieved returns near $9.8$). In the current asymmetric_advantages configuration with delay penalties the situation is qualitatively different. All baselines remain near the 40-40 expected penalty induced by the Delay wrapper and do not reliably escape into the high reward regime. Under any positive completion threshold, the completion rate is essentially zero for every method in every noisy regime.
To still obtain a notion of progress, we interpret small improvements relative to the 40-40 penalty floor as partial task completion. Let $R_{\min}=-40$ and define, for each $(m,e,r)$,
$C^{m,e,r}=\hat{V}^{m,e,r}-R_{\min}$ (21)
as the number of reward units recovered relative to the naive penalty baseline, and
$\tilde{C}^{m,e,r}=\frac{\hat{V}^{m,e,r}-R_{\min}}{0-R_{\min}}=\frac{\hat{V}^{m,e,r}+40}{40}$ (22)
as a normalized completion score that maps $R_{\min}$ to 0 and the clean zero reward benchmark to $1$. The task completion scripts compute seed aggregated quantities
$\bar{C}^{m,e}=\frac{1}{|\mathcal{S}_{\text{train}}|}\sum_{r}C^{m,e,r},\qquad\bar{\tilde{C}}^{m,e}=\frac{1}{|\mathcal{S}_{\text{train}}|}\sum_{r}\tilde{C}^{m,e,r},$
together with standard deviations across seeds.
(a) Clean regimes (No Noise and Noise).
(b) Perturbed regimes (Delay and Combo).
Figure 10: Task completion distributions under sparse rewards across clean and perturbed regimes (N = 5 seeds per configuration). In clean environments, all baselines achieve near-perfect normalized completion scores, indicating reliable coordination. Under delay and combined perturbations, completion performance degrades substantially across methods, highlighting the sensitivity of coordinated task execution to temporal and compounded disturbances.
For No Noise and Noise we have $\bar{V}^{m,e}=0$ for all $m$, so
$\bar{C}^{m,\text{No Noise}}=\bar{C}^{m,\text{Noise}}=40,\qquad\bar{\tilde{C}}^{m,\text{No Noise}}=\bar{\tilde{C}}^{m,\text{Noise}}=1,$
which simply reflects the fact that no delay penalty is applied. The more interesting behavior again appears in Delay and Combo, where Fig. 10(b) shows that all baselines lie extremely close to zero completion.
In the Delay regime the best methods only recover about one to one and a quarter reward units relative to the penalty floor:
TABLE V: Average coordination cost and normalized coordination cost under delay perturbations.
Method $\bar{C}^{(\cdot,\mathrm{Delay})}$ $\bar{\tilde{C}}^{(\cdot,\mathrm{Delay})}$
Baseline $\approx 1.26$ $\approx 0.0315$
CC_PPO $\approx 1.26$ $\approx 0.0315$
PBT_PPO $\approx 1.21$ $\approx 0.0303$
PPO+LLM $\approx 0.49$ $\approx 0.0123$
SP_PPO $\approx-0.13$ $\approx-0.0033$
HARL $\approx-0.18$ $\approx-0.0045$
Baseline and CC_PPO achieve the largest normalized completion scores at roughly $3.1\%$ of the full $40$ point gap, with PBT_PPO essentially tied. PPO+LLM recovers about $0.49$ reward units in Delay, or about $1.2\%$ of the maximum possible gap, which is lower than the best PPO variants but still strictly positive. HARL and SP_PPO slightly underperform the naive penalty baseline.
The Combo regime shows a different ranking:
TABLE VI: Average coordination cost and normalized coordination cost under combined noise perturbations.
Method $\bar{C}^{(\cdot,\mathrm{Combo})}$ $\bar{\tilde{C}}^{(\cdot,\mathrm{Combo})}$
Baseline $\approx-0.25$ $\approx-0.0063$
CC_PPO $\approx-0.25$ $\approx-0.0063$
PBT_PPO $\approx 0.83$ $\approx 0.0208$
PPO+LLM $\approx 0.81$ $\approx 0.0203$
SP_PPO $\approx 0.55$ $\approx 0.0138$
HARL $\approx-0.22$ $\approx-0.0055$
In words, PBT_PPO and PPO+LLM are tied within noise for the best normalized completion in Combo, recovering about $0.8$ reward units relative to the penalty floor and achieving completion scores around $2\%$ of the full gap. Baseline and CC_PPO fall slightly below the penalty floor on average, and HARL behaves similarly.
A one–way ANOVA on the normalized completion scores in Delay and Combo again yields no statistically significant differences across baselines (Delay: $F(5,24)\approx 1.31$, $p\approx 0.29$; Combo: $F(5,24)\approx 1.42$, $p\approx 0.25$). The main conclusion is that the asymmetric_advantages layout with delay penalties is extremely challenging under the current training budget. None of the methods consistently reach the high reward regime where full task completion becomes common; even the best methods recover only about one reward unit out of the $40$ point penalty. In this sense, the completion metric primarily highlights the difficulty of the environment rather than acting as a primary comparison signal. For this configuration, the combination of return statistics and Nash analysis provides a more informative view of the learned strategies.
III-E Equilibrium Behavior and Nash Gaps
Finally, we connect the episodic results with the Nash probe introduced in Section II-A. For each trained joint policy $\pi_{m,e,r}$ we compute its self play value $V_{\text{self}}^{m,e,r}$ and the value of a PPO based best response $V_{\text{BR}}^{m,e,r}$ trained in the best response environment $M_{\mathrm{BR}}(\pi_{m,e,r})$. The Nash gap
$\Delta^{m,e,r}=V_{\text{BR}}^{m,e,r}-V_{\text{self}}^{m,e,r}$ (23)
measures how much a unilateral deviation can improve the team return. Large positive gaps indicate exploitable policies; large negative gaps indicate that deviating hurts the deviating agent; gaps near zero correspond to approximate Nash behavior relative to the PPO best response class and training budget.
TABLE VII: Best response Nash analysis by regime. Values are averaged across all baselines and seeds.
Regime $\mathbb{E}[V_{\text{self}}]$ $\mathbb{E}[V_{\text{BR}}]$ $\mathbb{E}[\Delta]$
No Noise $0.00$ $0.00$ $0.00$
Noise $0.00$ $0.00$ $0.00$
Delay 40.33-40.33 39.77-39.77 $0.56$
Combo 39.62-39.62 40.35-40.35 0.73-0.73
(a) Clean regimes (No Noise and Noise).
(b) Perturbed regimes (Delay and Combo).
Figure 11: Distribution of Nash gaps $\Delta^{m,e,r}$ across clean and perturbed regimes (N = 5 seeds per configuration). In clean environments, Nash gaps remain tightly concentrated near zero across all baselines, indicating approximate equilibrium behavior under unperturbed dynamics. Under delay and combined perturbations, gaps increase substantially, reflecting heightened strategic instability and deviation from equilibrium across methods.
Aggregating the CSV produced by the best response analysis script yields three main observations.
  • In the No Noise and Noise regimes every model satisfies $\Delta^{m,e,r}=0$ up to numerical precision for all baselines, seeds, and joint policies. The best response agent is unable to improve or harm the team relative to self play, and the joint policies behave as fixed points of the PPO best response dynamics in these sparse settings.
  • In the Delay regime the Nash gaps are small but consistently positive. Averaging over all baselines and seeds we obtain
    $\mathbb{E}[\Delta^{\cdot,\text{Delay},\cdot}]\approx 0.57,\quad\operatorname{Std}[\Delta^{\cdot,\text{Delay},\cdot}]\approx 0.72,$
    with self play values and best response values
    $\mathbb{E}[V_{\text{self}}^{\cdot,\text{Delay},\cdot}]\approx-40.33,\qquad\mathbb{E}[V_{\text{BR}}^{\cdot,\text{Delay},\cdot}]\approx-39.77.$
    In $80\%$ of the $(m,r)$ combinations, the best response achieves a higher return than self play, but the magnitude is at most about two reward units. Relative to the 40-40 penalty floor, this corresponds to at most a five percent improvement, indicating mild exploitability rather than catastrophic instability.
  • In the Combo regime the gaps flip sign. The environment averaged value is
    $\mathbb{E}[\Delta^{\cdot,\text{Combo},\cdot}]\approx-0.73,\quad\operatorname{Std}[\Delta^{\cdot,\text{Combo},\cdot}]\approx 1.54,$
    with
    $\mathbb{E}[V_{\text{self}}^{\cdot,\text{Combo},\cdot}]\approx-39.62,\qquad\mathbb{E}[V_{\text{BR}}^{\cdot,\text{Combo},\cdot}]\approx-40.35.$
    Here the best response training procedure performs worse than the original joint policy, so unilateral deviations are on average harmful. In $80\%$ of the $(m,r)$ combinations we have $\Delta^{m,\text{Combo},r}<0$, which is consistent with joint strategies that have already learned to cope with both observation noise and delay penalties to the limited extent that the training budget allows.
Across all $120$ evaluated triples $(m,e,r)$, the distribution of Nash gaps is tightly concentrated. Half of the policies satisfy $|\Delta^{m,e,r}|\leq 0.25$, and $80\%$ satisfy $|\Delta^{m,e,r}|\leq 1$. Given that the absolute value of returns is on the order of $40$ in the challenging regimes, these small gaps indicate that all baselines, including PPO+LLM, behave as approximate Nash equilibria with respect to the PPO best response class under the given training horizon. Figure 11 shows that the vast majority of points lie in a narrow horizontal band around zero.